OpenStudy (anonymous):

How would i solve I x + 4 I= 3

OpenStudy (anonymous):

Can you tell me a few of attributes of absolute value?

OpenStudy (anonymous):

Thats all it gives me

OpenStudy (mertsj):

|x+4| =3 means that the number represented by x+4 is exactly 3 units from 0. The two numbers that are exactly 3 units from 0 are 3 and -3. Write and solve those two equations.

OpenStudy (mertsj):

\[x+4=3 \] or \[x+4=-3\]

OpenStudy (anonymous):

Oh wow was about to put the 3 and negative 3 in for X thanks

OpenStudy (anonymous):

You will be solving 2 equatons Absolute value means the distance of a point from 0. So it can be the positive or negative Hence your two equations will be x + 4 = 3 and x + 4 = -3 Solve both equations

OpenStudy (mertsj):

yw

OpenStudy (anonymous):

Thanks you all i think ive got it! How do i rate you guys up or whatever it is on this site?